OpenStudy (anonymous):

Solve and Check: 5(t + 1) + 8 = 3(t - 2) + 6

OpenStudy (anonymous):

5(t + 1) + 8 = 3(t - 2) + 6 5t+5+8=3t-6+6 5t-3t+13=0 2t=-13 t=-13/2=6.5

OpenStudy (anonymous):

Thank you so much! :)